Reflect on what you have learned about rational expressions and their applications, and then consider how you might apply rational expressions to your daily life. Explain this application, and discuss what the equation might be.
Answer by ankor@dixie-net.com(12702)

Here's a simple application on a tremendously important aspect of modern life
regarding pizza.
:
Let's say "Heartburn Pizza Shack" sells an 8" pizza for $5.00 and 16" pizza for
$12. Someone says, well heck, let's just get 2 ea 8" pizzas, that's a better deal.
:
But is it? The amt of pizza is determined by it's area
The 8" pizza (r=4"):
A = pi*4^2
A = 50.265 sq in. Two of them would be 100.53 sq in
:
Cost 10/100.53 = .099 or let's say 10 cents per sq inch
:
The 16" pizza (r=8)
A = pi*8^2
A = 201 sq inches (1 ea 16" pizza)
:
Cost 12/201 = .059 or about 6 cents per sq inch
:
Hope this qualifies as real life. I had fun with this with my grand-kids while
waiting to be served in a local pizza place recently.